# Auto-Compressing Networks

### v1

@@ 3 ACNs in Practice: Information Compression and Gradient Flow
In this section, we move from theoretical analysis and simple demonstrations to implementing auto-compressing networks in modern deep learning architectures. We apply our approach to state-of-the-art neural network models across diverse tasks and datasets, demonstrating that the information compression effect observed in our theoretical analysis manifests consistently in practice.
Experimental Setup:
Our experimental validation spans multiple domains and model architectures. We implement ACNs using variants of the Transformer (Vaswani, 2017) for language and vision tasks and MLP-Mixer (Tolstikhin et al., 2021) for vision tasks. This allows us to evaluate our approach on diverse benchmarks including image classification (CIFAR-10, ImageNet-1K), sentiment analysis, and language understanding. In each implementation, we follow the core ACN design principle: for each input token, we compute a final output vector $y^{t}$ (where $t$ is the sequence index) by summing the output representations of all intermediate layers along with the input embedding, as shown in Eq. 1. To generate classification predictions, we either apply a pooling layer to these vectors for image classification or use the final representation of the [class] ([CLS]) token for text classification. For a network of depth $L$, making predictions using $k$ intermediate layers involves computing $y_{k}^{t}$ for each token, which is the sum of intermediate representations up to layer $k$. This summed representation is then passed to a single global classification head, which is trained once and shared across all sub-networks (we do not retrain or create separate classification heads for each depth configuration). This approach yields $L+1$ sub-networks, ranging from using only the input embedding (the first sub-network) to the full network (the last sub-network). For example, the network shown in Figure 11(c) would be the $L-2$ subnetwork (utilizing layers $1$ to $L-2$), whereas the network in in Figure 11(b) would be the full network (all layers included). When evaluating residual network baselines, we follow standard practice: to assess the network at depth $k$, we simply take the output $y_{k}^{t}$ of the $k$th layer as our representation. This provides a natural comparison point to ACNs at equivalent depths. All other procedures remain the same. In all figures, prediction layer 0 refers to the input embedding passed through the classification head for prediction. Additional experimental details and hyperparameters are provided in Appendix A.
3.1 Auto-Compression via Direct Gradient Flow and Layer-wise training
@@
Figure 3: (left) ACNs is the only architectural variant that achieve auto-compression. (right) The ratio of direct gradient DG to the total gradient FG in auto-compressing vs residual architectures. The exponential (vs linear) number of paths in Residual Networks decreases the influence of DG in the training dynamics of the network compared to Auto-Compressing Networks.
@@
4 Compression Capabilities of Auto-Compressing Networks
4.1 Auto-Compressing Vision Transformers
@@
Next, we evaluate ACNs in the context of transformer architectures by implementing an auto-compressing variant of Vision Transformer (ViT) (Dosovitskiy, 2020). We train a Vision Transformer (ViT) with long connections (AC-ViT) from scratch on the ILSVRC-2012 ImageNet-1K, following the training setup in the original paper. For both models we use $256$ batch size due to memory constraints. AC-ViT converges at $700$ epochs, while the Residual ViT converges at $300$ epochs [^10] In this more challenging setting, we observe a trade-off between training and inference time, which is partially aleviated using a parameterization similar to DiracNets (Zagoruyko & Komodakis, 2017) for the MLP layers, specifically $\hat{W}=(I+W)$.. As shown in Fig. 4, AC-ViT reaches top performance at only $6$ layers while the vanilla ViT needs all $12$ layers to reach similar performance, effectively suggesting that ACNs can improve inference time and memory consumption without sacrificing performance. To gain more intuition about the training dynamics and task learning of the two variants, in Figure 1(right) we plot the incremental layer performance contribution (difference in accuracy of subnetwork $i+1$ to subnetwork $i$) to track the behavior of intermediate layers throughout training. The key observation is that ACNs (right) are trained in a layer-wise fashion where early layers are trained at a faster rate and task-relevant information is gradually pushed only to a subset of the deeper layers, achieving strong performance along with auto-compression. In the contrary, the Residual variant performs task-learning in the last 2-3 layers, effectively utilizing the full network to achieve top performance.
4.2 The Effect of Task Difficulty
Intuitively, overparameterized networks trained on easier tasks should demonstrate higher levels of redundancy. Therefore, ACNs should converge to utilizing fewer layers as task difficulty decreases. To verify this, we use the number of classes as a proxy for task difficulty for image classification on the CIFAR-10 dataset (Krizhevsky, 2009). Specifically, we create subsets of $2$, $5$, and $10$ classes, the assumption being that binary classification should be easier than $10$-class classification. For this experiment we utilize MLP-Mixer (Tolstikhin et al., 2021) and train two variants, the original MLP-Mixer with residual connections and the modified MLP-Mixer with long connections (AC-Mixer). Results are presented in Fig 5. We observe that indeed AC-Mixer converges to solutions with larger effective depth, as the task “difficulty” increases. Specifically, in this experiment, ACN needs $8$, $10$ and $12$ layers for the $2$, $5$ and $10$-class classification problem, respectively. In contrast, the Residual Mixer converges to solutions where the full depth of the network is utilized, irrespective of the task difficulty  [^11] The Residual Mixer was trained for 300 epochs, while AC-Mixer for 420 epochs to reach the performance of its residual counterpart..
@@ 4.2 The Effect of Task Difficulty
Figure 5: Performance of the intermediate layers as the number of classes (and examples) in the CIFAR-10 dataset increases from 2, to 5 to 10 classes: (a) Residual Mixer vs (b) AC-mixer.
5 Generalization Capabilities of Auto-Compressing Networks
@@
While ACNs demonstrate effective parameter reduction through architectural compression, a key question remains: do these compressed representations offer additional benefits beyond parameter efficiency? In this section, we investigate whether the concentrated information in ACNs’ early layers leads to improved generalization capabilities compared to traditional residual architectures. Specifically, we explore two critical aspects of generalization: robustness to input noise and performance in low-data regimes.
@@ 5.1 Robustness to Input Noise
Next, we present results assessing the robustness of ACNs versus residual transformer architectures to input noise. The experiments are performed with the AC-ViT and residual ViT architectures trained on ImageNet-1K. In this experiment, we inject increasing levels of additive Gaussian noise with standard deviation $\sigma=0.1,0.2,0.4$, and salt-and-pepper noise with percentage of altered pixels $p=1\%,2\%,10\%$. Results (average accuracy) are shown in Table 2 (a) for Gaussian and (b) for salt-and-pepper noise. We observe that ACNs display improved robustness to noise, and the performance gap with the residual transformer increases as the noise levels increase. These results align with the findings of Yang et al. (2020), who showed that architectures with forward passes closer to feedforward networks (like our ACNs) exhibit enhanced noise robustness. In residual architectures, short connections allow noise to propagate and accumulate throughout the network, whereas the long-connection design of ACNs helps mitigate this amplification effect.
@@ 5.2 Robustness to Data Sparsity
Next, we experimentally compare the performance of residual and long connections architectures in low-data scenarios. For this purpose, we create a random subset of CIFAR-10 (Krizhevsky, 2009) by retaining only $100$ samples per class, resulting in a total of $1000$ examples. Using the same training settings and models as described in Section 4.2, we train both architectures for $150$ epochs to assess how fast the training and test loss decrease, as a proxy for the generalization capabilities of each architecture. Results shown in Fig. 6 reveal that ACNs achieve lower training and test loss in fewer epochs compared to residual networks. This faster convergence in loss metrics is a strong indication that auto-compressing networks can be effectively utilized in scenarios with limited data.
6 Auto-Compressing Encoder Architectures for Language Modeling
Recent studies have demonstrated substantial parameter redundancy in modern foundation models, particularly in their deeper layers (eg. Gromov et al. (2024)). This characteristic is crucial today, in the context of large language and multimodal models, which are typically pre-trained as general-purpose models before being adapted to specific downstream tasks. Since these specialized applications may not require the full parameter capacity of the base model, learned representations (through architectural choices) that facilitate subsequent compression and pruning become crucial. In this section, we conduct a preliminary study on the effectiveness of the ACN architecture in general pre-training (masked language modeling with a BERT architecture) followed by fine-tuning and pruning. The results show that ACNs learn compact representations that: 1) achieve on-par performance with the residual architecture on transfer learning tasks, while utilizing significantly fewer parameters, and 2) complement post-training pruning techniques, enhancing their effectiveness.
6.1 Masked Language Modeling and Transfer Learning with ACNs
Next, we compare the ACN and residual architectures in the standard BERT pre-training and fine-tuning paradigm. Using the original BERT pretraining corpus (BooksCorpus (Zhu et al., 2015) and English Wikipedia), we train both architectures to equivalent loss values; the AC-BERT variant requires two epochs vs one epoch for the residual baseline. Following pre-training, we fine-tune both models on three GLUE benchmark datasets (Wang et al., 2018a): SST-2 sentiment analysis (Socher et al., 2013), QQP paraphrasing, and QNLI question answering (Rajpurkar et al., 2016).
Figure 7(left) demonstrates a key advantage of the ACN architecture: it naturally converges to using significantly fewer layers (approximately 75% less layers) while maintaining performance comparable to the full residual network. These results suggest promising applications for ACNs in large language models, where pre-training could be performed with long connections, allowing downstream tasks to adaptively utilize only the necessary subset of layers during fine-tuning.
Figure 7: (left) Downstream performance of AC-BERT vs residual BERT on three GLUE tasks: sentiment analysis (SST-2), paraphrasing (QQP), and question answering (QNLI). (right) Accuracy vs model size of AC-BERT and Residual BERT on SST-2 when pruned with Magnitude and Movement Pruning (with two different settings, refer to Appendix A for details).
6.2 Post-Training Pruning with AC-Encoders
ACN’s primary advantage lies in its inherent compression capabilities during training, suggesting that when combined with pruning techniques, it should significantly outperform traditional residual architectures. To provide validation for this hypothesis, we conducted experiments using magnitude and movement pruning (Sanh et al., 2020), two commonly employed baseline pruning techniques. Results are shown when fine-tuning of the SST-2 dataset sentiment analysis task. We refer to Appendix A for details regarding the experimental setup. Figure 7(right) confirms our hypothesis: ACNs consistently demonstrate superior compression-performance trade-offs compared to standard architectures, with their advantage becoming more pronounced at higher compression rates. This indicates that ACNs’ architectural design naturally leads to more efficient parameter utilization, creating representations that are inherently more amenable to further pruning. While these preliminary results validate our approach to addressing parameter redundancy, they also point toward promising future directions. We anticipate that combining pre-trained ACN architectures with state-of-the-art pruning methods will result in extremely efficient, high-performing models, though rigorous validation of this hypothesis requires further investigation.
7 Mitigating Catastrophic Forgetting with ACNs
@@
Continual learning involves training models on a sequence of tasks without access to past data, aiming to retain performance on previous tasks while learning new ones (De Lange et al., 2021; Wang et al., 2024). A central challenge in CL is catastrophic forgetting—the tendency of neural networks to overwrite old knowledge when updated with new data. Common approaches include data replay methods (Rebuffi et al., 2017; Lopez-Paz & Ranzato, 2017) and regularization techniques that penalize changes to important parameters (Kirkpatrick et al., 2017; Zenke et al., 2017; Aljundi et al., 2018). We’ve already demonstrated that ACNs, through implicit layer-wise training, dynamically allocate parameters based on task demands while preserving redundant parameters for future tasks. Conversely, Residual Networks optimized for efficient task learning risk overfitting and suboptimal parameter usage in these sequential learning settings. To test our claims, we evaluate both architectures on the split CIFAR-100 continual learning benchmark, comprising 20 sequential disjoint 5-class classification tasks, focusing on task-incremental learning (Van de Ven & Tolias, 2019) where task identity is known. We utilize MLP-Mixer architectures (hyperparameters in Appendix A) and we test two continual learning algorithms trained for 10 epochs for each task: naive fine-tuning (Naive FT) and Synaptic Intelligence (SI) (Zenke et al., 2017), which adds a gradient-based regularizer to each parameter depending on how changes in it affect the total loss in a task over the training trajectory. Across experiments, we report Average Forgetting, defined as the mean difference between a task’s best performance (right after it is learned) and its final performance after all tasks are learned, and Average Accuracy, defined as the mean accuracy over all tasks at the end of training. We expect gradient-based regularization methods to perform particularly well with ACNs since unused, redundant parameters receive small gradients, making their detection easier compared to Residual Networks where gradients are more uniformly distributed (see gradient heatmaps, Fig. 1(left)). Results in Table 3 confirm our intuition: ACNs consistently exhibit significantly less forgetting (up to 18% improvement) compared to Residual Networks. Notably, with SI, increasing ACN depth decreases forgetting—an ideal behavior for CL systems where increasing network capacity reduces forgetting—while Residual Networks show the opposite pattern, indicating potential overfitting. ACNs also achieve better average accuracy across all tasks, further establishing them as a more suitable architecture for continual learning.
@@
8 Auto-Compressing Architectures vs. Layer-wise Loss Regularization
@@
Parameter redundancy, and specifically potential layer redundancy, in residual architectures is a phenomenon that has been well documented (Alain & Bengio, 2016; Veit et al., 2016; Huang et al., 2016). Recent works (Elhoushi et al., 2024b; Jiang et al., 2024) have attempted to address this through regularization-based layer-wise structural learning approaches during training, specifically by adding losses to all intermediate layers of the network and using a weighted sum of them as the total loss, a technique formally introduced in (Lee et al., 2015) for improved training. Such loss-based regularization methods rely heavily on precise tuning of intermediate loss weights, creating practical challenges. If early-layer loss weights are set too high, the network risks overfitting and poor generalization; if set too low, performance improves gradually across layers with no clear cutoff point, reaching optimal results only at the final layer. This sensitivity to hyperparameter selection makes it difficult to reliably identify an optimal depth for inference using loss-based regularization. ACNs address this challenge through architectural design rather than regularization, naturally compressing information without requiring complex hyperparameter tuning.
8.1 ACNs achieve better generalization
To evaluate hyperparameter sensitivity in regularization-based approaches, we compare several methods on the CIFAR-10 dataset using MLP-Mixer architectures. Our comparison includes: 1) our proposed AC-Mixer, 2) an unregularized Residual Mixer as baseline, 3) a Residual Mixer with the setup of (Jiang et al., 2024) (Aligned), 4) a Residual Mixer with the setup of (Elhoushi et al., 2024b) (LayerSkip), with the rotational early exit curriculum with $p_{max}=0.1$, $e_{scale}=0.2$ and $C_{rot,R}=15$, 5) a Residual Mixer with a baseline vanilla deep supervision (Lee et al., 2015) where all intermediate losses before the final layer are weighted with $\lambda=0.1$ (DeepSup). This comparative analysis reveals how different approaches respond to their respective hyperparameter configurations. We follow the training pipeline as described in the previous section and track the performance of all intermediate layers. Our experiments (Figure 8) demonstrate that while regularization approaches are highly sensitive to intermediate loss weights, creating a trade-off between performance and compression, ACNs consistently achieve strong results through their inherent architectural properties. Specifically, ACNs match the performance of unregularized Residual Networks, while effectively determining a shallower cutoff layer. While careful tuning of regularization methods can potentially match ACN’s performance, our approach provides a more elegant and robust solution that requires no parameter adjustment while maintaining high performance and achieving sparsity.
Figure 8: Performance of intermediate layers on CIFAR-10 of AC-Mixer, unregularized Residual Mixer and Residual-Mixer with intermediate losses based on (Jiang et al., 2024) (Aligned), (Elhoushi et al., 2024b) (LayerSkip) and (Lee et al., 2015) (DeepSup). For better resolution only the last 4 layers are shown in the plot.
8.2 ACNs show stronger Transfer Learning capabilities

### v2

@@ 3 ACNs in Practice: Information Compression and Gradient Flow
Next, we implement auto-compressing networks on top of state-of-the-art neural architectures across diverse tasks and datasets. We implement ACNs using variants of the Transformer (vaswani2017attention, ) for language and vision tasks and MLP-Mixer (tolstikhin2021mlp, ) for vision tasks. This allows us to evaluate our approach on diverse benchmarks including image classification (CIFAR-10 (krizhevsky2009learning, ), ImageNet-1K (russakovsky2015imagenet, )), sentiment analysis, and language understanding (BERT (devlin2019bert, ) on GLUE (wang-etal-2018-glue, )).
In ACNs, for each input token, the final output vector $y^{t}$ is the sum of all intermediate layer representations plus the input embedding (Eq. 1). For classification, we apply pooling for images or use the [CLS] token for text. For a network of depth $L$, predictions using $k$ intermediate layers compute $y_{k}^{t}$ (sum of representations up to layer $k$) for each token, passed to a single shared classification head. This yields LL+1 sub-networks, from input-only to full-network. For residual network baselines, at depth $k$, we take the output $y_{k}^{t}$ of layer $k$ as our $k+1$ subnetwork’s output.
@@
Figure 2: Results for the CIFAR-10 task using the MLM-Mixer base architecture: (left) ACN variants achieve auto-compression needing only a few layers to achieve good performance. (right) The ratio of direct gradient DG to the total gradient FG is significantly higher in early layers for ACNs.
@@
4 ACNs compress more
4.1 Auto-Compressing Vision Transformers and MLP-Mixers
@@
Next, we evaluate ACNs in the context of transformer architectures by implementing an auto-compressing variant of Vision Transformer (ViT) (dosovitskiy2021an, ). We train a Vision Transformer (ViT) with long connections (AC-ViT) from scratch on the ILSVRC-2012 ImageNet-1K, following the training setup in the original paper. For both models we use $256$ batch size due to memory constraints. AC-ViT converges at $700$ epochs, while the Residual ViT converges at $300$ epochs. As shown in Fig. 3, AC-ViT reaches top performance at only $6$ layers while the vanilla ViT needs all $12$ layers to reach similar performance, effectively suggesting that ACNs can improve inference time and memory consumption without sacrificing performance. To gain more intuition about the training dynamics and task learning of the two variants, in Figure 1(right) we plot the incremental layer performance contribution (difference in accuracy of subnetwork $i+1$ to subnetwork $i$) to track the behavior of intermediate layers throughout training. The key observation is that ACNs are trained in a layer-wise fashion, while the residual variant performs task-learning in the last 2-3 layers, effectively utilizing the full network to achieve top performance.
@@
Figure 4: Performance of the intermediate layers as the number of classes (and examples) in the CIFAR-10 dataset increases from 2, to 5 to 10 classes: (a) Residual Mixer vs (b) AC-mixer (C denotes the number of classes in the subset).
The effect of Task Difficulty
Intuitively, overparameterized networks trained on easier tasks should demonstrate higher levels of redundancy. Therefore, ACNs should converge to utilizing fewer layers as task difficulty decreases. To verify this, we use the number of classes as a proxy for task difficulty for image classification on the CIFAR-10 dataset (krizhevsky2009learning, ). Specifically, we create subsets of $2$, 55, and $10$ classes, the assumption being that binary classification should be easier than $10$-class classification. For this experiment we utilize MLP-Mixer (tolstikhin2021mlp, ) and train two variants, the original MLP-Mixer with residual connections and the modified MLP-Mixer with long connections (AC-Mixer). Results are presented in Fig 4. We observe that indeed AC-Mixer converges to solutions with larger effective depth, as the task “difficulty” increases. Specifically, in this experiment, ACN needs $8$, $10$ and $12$ layers for the $2$, 55 and $10$-class classification problem, respectively. In contrast, the Residual Mixer converges to solutions where the full depth of the network is utilized, irrespective of the task difficulty  [^6] The Residual Mixer was trained for 300 epochs, while AC-Mixer for 420 epochs to reach the performance of its residual counterpart..
4.2 Auto-Compressing Encoder Architectures for Language Modeling
In this section, we conduct a preliminary study on the effectiveness of the ACN architecture in general pre-training (masked language modeling with a BERT architecture) followed by fine-tuning. The results show that ACNs learn compact representations that: 1) achieve on-par performance with the residual architecture on transfer learning tasks, while utilizing significantly fewer parameters, and 2) complement post-training pruning techniques, enhancing their effectiveness.
4.2.1 Masked Language Modeling and Transfer Learning with ACNs
We compare the ACN and residual architectures in the standard BERT pre-training and fine-tuning paradigm. Using the original BERT pretraining corpus (BooksCorpus (zhu2015aligning, ) and English Wikipedia), we train both architectures to equivalent loss values; the AC-BERT variant requires two epochs vs one epoch for the residual baseline. Following pre-training, we fine-tune both models on three GLUE benchmark datasets (wang-etal-2018-glue, ): SST-2 sentiment analysis (socher-etal-2013-recursive, ), QQP paraphrasing, and QNLI question answering (rajpurkar-etal-2016-squad, ). Figure 5(left) demonstrates a key advantage of the ACN architecture: it naturally converges to using significantly fewer layers (approximately 75% less layers) while maintaining performance comparable to the full residual network. These results suggest promising applications for ACNs in large language models, where pre-training could be performed with long connections, allowing downstream tasks to adaptively utilize only the necessary subset of layers during fine-tuning. To further enhance compression, ACNs can be combined with standard pruning techniques—a preliminary investigation of this approach is provided in Appendix G.
Figure 5: Downstream performance of AC-BERT vs residual BERT on three GLUE tasks.
5 ACNs generalize better
@@
While ACNs demonstrate effective parameter reduction through architectural compression, a key question remains: do these compressed representations offer additional benefits beyond parameter efficiency? In this section, we investigate whether the concentrated information in ACNs’ early layers leads to improved generalization capabilities compared to traditional residual architectures. Specifically, we explore robustness to input noise as a proxy for generalization. Moreover, we compare the inherent auto-compression of ACNs—achieved through architectural design—to recent regularization-based methods that rely on externally imposed intermediate losses.
@@ 5.1 Robustness to Input Noise
Next, we present results assessing the robustness of ACNs versus residual transformer architectures to input noise. The experiments are performed with the AC-ViT and residual ViT architectures trained on ImageNet-1K. In this experiment, we inject increasing levels of additive Gaussian noise with standard deviation $\sigma=0.1,0.2,0.4$, and salt-and-pepper noise with percentage of altered pixels $p=1\%,2\%,10\%$. Results (average accuracy) are shown in Table 2 (a) for Gaussian and (b) for salt-and-pepper noise. We observe that ACNs display improved robustness to noise, and the performance gap with the residual transformer increases as the noise levels increase. These results align with the findings of (yang2020interpolation, ), who showed that architectures with forward passes closer to feedforward networks (like our ACNs) exhibit enhanced noise robustness. In residual architectures, short connections allow noise to propagate and accumulate throughout the network, whereas the long-connection design of ACNs helps mitigate this amplification effect.
@@ 5.2 Robustness to Data Sparsity
Next, we experimentally compare the performance of residual and long connections architectures in low-data scenarios. For this purpose, we create a random subset of CIFAR-10 (krizhevsky2009learning, ) by retaining only $100$ samples per class, resulting in a total of $1000$ examples. Using the same training settings and models as described in Section 4.1 (MLP-Mixer on CIFAR-10), we train both architectures for $150$ epochs to assess how fast the training and test loss decrease, as a proxy for the generalization capabilities of each architecture. Results shown in Fig. 6 reveal that ACNs achieve lower training and test loss in fewer epochs compared to residual networks. This faster convergence in loss metrics is a strong indication that auto-compressing networks can be effectively utilized in scenarios with limited data.
6 ACNs forget less
@@
Continual learning involves training models on a sequence of tasks without access to past data, aiming to retain performance on previous tasks while learning new ones (de2021continual, ; wang2024comprehensive, ). A central challenge in CL is catastrophic forgetting—the tendency of neural networks to overwrite old knowledge when updated with new data. Common approaches include data replay methods (rebuffi2017icarl, ; lopez2017gradient, ) and regularization techniques that penalize changes to important parameters (kirkpatrick2017overcoming, ; zenke2017continual, ; aljundi2018memory, ). We’ve already demonstrated that ACNs, through implicit layer-wise training, dynamically allocate parameters based on task demands while preserving redundant parameters for future tasks. Conversely, Residual Networks optimized for efficient task learning risk overfitting and suboptimal parameter usage in these sequential learning settings. To test our claims, we evaluate both architectures on the split CIFAR-100 continual learning benchmark, comprising 20 sequential disjoint 5-class classification tasks, focusing on task-incremental learning (van2019three, ) where task identity is known. We utilize MLP-Mixer architectures (hyperparameters in Appendix B) and we test two continual learning algorithms trained for 10 epochs for each task: naive fine-tuning (Naive FT) and Synaptic Intelligence (SI) (zenke2017continual, ), which adds a gradient-based regularizer to each parameter depending on how changes in it affect the total loss in a task over the training trajectory. Across experiments, we report Average Forgetting, defined as the mean difference between a task’s best performance (right after it is learned) and its final performance after all tasks are learned, and Average Accuracy, defined as the mean accuracy over all tasks at the end of training. We expect gradient-based regularization methods to perform particularly well with ACNs since unused, redundant parameters receive small gradients, making their detection easier compared to Residual Networks where gradients are more uniformly distributed (see gradient heatmaps, Fig. 1(left)). Results in Table 3 confirm our intuition: ACNs consistently exhibit significantly less forgetting (up to 18% improvement) compared to Residual Networks. Notably, with SI, increasing ACN depth decreases forgetting—an ideal behavior for CL systems where increasing network capacity reduces forgetting—while Residual Networks show the opposite pattern, indicating potential overfitting. ACNs also achieve better average accuracy across all tasks, further establishing them as a more suitable architecture for continual learning.
@@
7 ACNs transfer better
@@
Parameter redundancy, and specifically potential layer redundancy, in residual architectures is a phenomenon that has been well documented (alain2016understanding, ; veit2016residual, ; huang2016deep, ). Recent works (elhoushi2024layerskip, ; jiang2025tracing, ) have attempted to address this through regularization-based layer-wise compression techniques during training, specifically by adding losses to all intermediate layers of the network and using a weighted sum of them as the total loss, a technique formally introduced in (lee2015deeply, ) for improved training. Such loss-based regularization methods rely heavily on precise tuning of intermediate loss weights, creating practical challenges. If early-layer loss weights are set too high, the network risks overfitting and poor generalization; if set too low, performance improves gradually across layers with no clear cutoff point, reaching optimal results only at the final layer. This sensitivity to hyperparameter selection makes it difficult to reliably identify an optimal depth for inference using loss-based regularization. ACNs address this challenge through architectural design rather than regularization, naturally compressing information without requiring complex hyperparameter tuning.
Method Accuracy (%)
AC-Mixer 85.38 $\pm$0.7
Aligned 82.9 $\pm$0.9
LayerSkip 79 $\pm$1.2
Table 4: C-100 to C-10 transfer learning performance of ACNs vs recent regularization-based layer copression approaches.